OpenStudy (anonymous):

If f(x) = |x – 5| + 2, find f(3). @gigi.antonia

OpenStudy (anonymous):

\(f(3)=|(3)-5|+2\)

OpenStudy (anonymous):

4

OpenStudy (anonymous):

So does that make it 4?

OpenStudy (anonymous):

So once you plug it in solve within the absolute value sign and remember it is always positive. At the end, add 2 to your answer.

OpenStudy (anonymous):

\(f(3)=|3-5|+2=|-2|+2=2+2=4\)

OpenStudy (anonymous):

ok that helps a bit but I'm still not confident with the whole process.

OpenStudy (anonymous):

3-5=-2 , but the absolute value signs make it 2, then add the ending, 2+2=4

OpenStudy (anonymous):

Could you maybe help me on another?

OpenStudy (anonymous):

Of course (:

OpenStudy (anonymous):

absolute value is just distance. i live in san diego and tia juana is 30 miles south and la is 100 miles north. they are in opposite directions but i don't say tj is -30 miles and la is +100. they're 30 and 100, resp., but in different directions. that's absolute value. it's just the distance, not the direction.